### 15 Mar 2024 v4

# Jordan Algebraic Formulation of Quantum Mechanics and The Non-commutative Landau Problem

Jordan Algebraic Formulation of Quantum Mechanics
and
The Non-commutative Landau Problem
[…]
( ${}^{1}$ Department of Physics, College of Sciences, Koç University, 34450 Sarıyer, Istanbul, Turkey
${}^{2}$ Faculty of Engineering and Natural Sciences, Maltepe University, 34857 Maltepe, Istanbul,Turkey
March 15, 2024)
We present a Jordan algebraic, non-associative formulation of the non-commutative Landau problem coupled to a harmonic potential. To achieve this, an alternative formulation of the Hilbert space version of quantum mechanics is presented. Using this construction, the Hilbert space corresponding to the non-commutative Landau problem is obtained. Non-commutative parameters are then described in terms of an associator in the Jordan algebraic setting. Pure states and density matrices arising from this problem are characterized. This in turn leads us to the Jordan-Schrödinger time-evolution equation for the state vectors for this specific problem.
1 Introduction
In the algebraic formulations of quantum mechanics, observables play a central role. The main motivation for such approaches is to formulate quantum mechanics in terms of essential ingredients only; which are the observables, states, expectation values and their time evolution. For instance, it is possible [1] to take quantum mechanical observables as the self-adjoint part of a $C^{*}$-algebra $\mathcal{A}$ and states as positive linear functionals on $\mathcal{A}$. Then the real number that a state assigns to an observable will be the expectation value of that observable in that state [^3] See [2], [3], [4] for deeper formulations of quantum mechanics.. However, it turns out that the $C^{*}$-algebra one uses in such approaches usually allows for non-self-adjoint entities. Consequently, to start with, a more restrictive structure is necessary to have a consistent description of the algebra of observables.
The Jordan algebraic formulation [5] adopted here is based on the fact that quantum mechanical observables form an algebra $\mathcal{A}$ under the non-associative Jordan product:
$A\bullet B=\frac{1}{2}(AB+BA).$
This algebra qualifies as a real Jordan algebra of operators since the Jordan product is commutative and satisfies the Jordan identity
$(A\bullet B)\bullet A^{2}=A\bullet(B\bullet A^{2}).$
This last relation is equivalent to power-associativity of $\mathcal{A}$ when $\mathcal{A}$ is formally real [5]. Furthermore, a Jordan algebra is called special if it comes from the Jordan product in an associative algebra, which is the case in present definition of Jordan algebra of observables. Since the algebras arising in quantum mechanics are typically infinite dimensional, it is of interest to develop a structure theory compatible with the infinite dimensional case. This idea is initiated by [6], where the direct analogues of $C^{*}$-algebras that are called $JB$-algebras are studied. $JB$-algebras are Jordan algebras with a Banach space structure. They satisfy the following conditions:
$\|A^{2}\|=\|A\|^{2},$
$\|A^{2}-B^{2}\|\leq\max\{\|A^{2}\|,\|B\|^{2}\},$
$\|A\bullet B\|\leq\|A\|\|B\|.$
The first condition, also known as the $C^{*}$ condition, is a consequence of the second (positivity condition) and the third conditions (an analogue of Banach space axiom). Therefore, the observables, which can be considered as the self-adjoint part of a $C^{*}$-algebra, form a $JC$-algebra, or in other words, a special $JB$-algebra. At this point, we can say that if one wants to formulate quantum mechanical problems using the essential ingredients only, it is important to write down a formulation using the Jordan algebraic approach. Also, we note that all $JB$-algebras, with the exception of $H_{3}(\mathbb{O})$, can be seen as the self-adjoint part of a $C^{*}$-algebra, so they can be represented on a complex Hilbert space [6]. We refer to [7], [8], [9], [10] for more detailed work on the exceptional character of $H_{3}(\mathbb{O})$, and Jordan algebras.
We would like to note that non-associative structures found applications in physical problems [11], [12], [13], [14], [15], [16]. The associative product used in quantum mechanics do not apply to some exotic systems such as magnetic monopoles, in such problems a non-associative extension of quantum mechanics is vital. It turns out that if we consider dynamics of electrons in uniform distributions of magnetic charge, certain non-associative deformations of quantum mechanics and gravity in three dimensions are needed. The quantitative framework for non-associative quantum mechanics in this setting exhibits new effects compared to ordinary quantum mechanics with sourceless magnetic fields. On the other hand, the non-associativity of translations in a quantum system with a background magnetic field has received renewed interest in association with topologically trivial gerbes over $\mathbb{R}^{n}$. Then the non-associativity arises when trying to lift the translation group action on the 1-particle system to a second quantized system, described by a 3-cocycle of the group $\mathbb{R}^{n}$ with values in the unit circle $S^{1}$.
In this paper, our main interest will be the Landau problem that involves the motion of an electron interacting with a constant, uniform background magnetic field. The minimal coupling of the electron charge density to an external magnetic field is described in terms of the electromagnetic vector potential $\mathbfcal{A}$ and leads to the Hamiltonian in the symmetric gauge
$H=\frac{1}{2m}\left({\bf\it p}-\frac{e}{c}\mathbfcal{A}\right)^{2}=:\frac{1}{2m}{\bf\it P}^{2},$
where ${\bf\it P}$ denotes kinetic momenta. Thus, the Landau problem will relate to the quantum mechanics of a charged particle on the transverse $xy$-plane in the presence of a constant, uniform magnetic field present along the $z$-axis. In metals, electrons occupy Landau energy levels given by
$E_{n}=\hbar\omega\left(N+\frac{1}{2}\right),$
where the Larmor frequency $\omega_{L}=\frac{eB}{mc}$ and $N=n+\frac{m-|m|}{2}$. The infinitely degenerate Landau levels are labeled by two quantum numbers; the radial quantum number $n\geq 0$, and the magnetic quantum number $m$ with range $m\geq-n$. The energy spectrum is bounded from below: $N\geq 0$.
From the physical point of view, the Landau problem turned out to be one of the most engaging problems in quantum mechanics. It is inspired by ideas in diverse areas of physics such as quantum optics and condensed matter physics. Since it finds numerous applications in different areas of physics, the Landau problem is still being studied extensively in the literature. In particular, the first author et al. [17] showed recently that a Newton-Hooke duality exists between the $2D$ Hydrogen atom and the Landau problems via the Tits-Kantor-Koecher construction of the conformal symmetries of the Jordan algebra of real symmetric $2\times 2$ matrices. Here we seek a further application of the Jordan algebraic approach to the non-commutative Landau problem.
Non-commuting spatial coordinates and fields can be seen in actual physical situations [18]. Generalizations of canonical commutation relations in the traditional approach lead to non-commutative geometry and appear in quantum mechanics with source-free magnetic fields. This generalization of the Landau problem to the case of non-commutative quantum mechanics has been actively studied in the context of physics associated with non-commutative geometry [19], [20], [21], [22], [23], [24], [25], [26], [27], [28], [29], [30], [31], [32], [33], [34]. In [19], a correspondence is established between the dynamics of the two-vortex system and the non-commutative Landau problem. For discussions of the exotic Galilean symmetry of a particle in the non-commutative plane, see [20], [21], [22], [29]. One of the main ingredients of this paper, namely the chiral decomposition in the non-commutative Landau problem, is advocated in [24], [25] and is further generalized to nonvanishing electric fields in [34]. In this paper, different from the previous works, we prefer to use a non-associative approach to handle the non-commutative Landau problem. It turns out that the traditional (associative) approach is incapable of handling generalizations of the non-commutative Landau problem, namely the problems containing magnetic charges. But our approach here has the advantage of being suitable for such generalizations.
Given an electron moving in a plane in the presence of a constant external magnetic field ${\bf\it B}$ perpendicular to that plane, the spectrum of the quantized theory gives infinitely degenerate Landau levels, with separation $\mathcal{O}\left(\frac{eB}{mc}\right)$. The limit $B\rightarrow\infty$ or $m\rightarrow 0$ projects onto the lowest Landau level. On each of the projected Landau level, one obtains a non-commuting algebra for the space coordinates giving the center of the cyclotron motion,
$[X,Y]=-\frac{i\hbar c}{eB}\neq 0.$
The main motivation of this paper follows from the proof below that the non-commutativity of spatial coordinates that involve only individual Landau levels can be cast as non-associativity of observables based on a corresponding Jordan algebraic formulation.
In what follows, we formulate the non-commutative Landau problem coupled to a harmonic potential in the Jordan algebraic approach we give below. It turns out that, in this problem, one obtains a pair of commuting Hamiltonians, which can be written in terms of two pairs of mutually commuting creation and annihilation operators of oscillator type. These operators further generate two mutually commuting von Neumann algebras. Based on the Hilbert-Schmidt operator formulation of the non-commutative Landau problem, it is possible to write down Jordan algebraic formulas. At this point, we take the associative case as our guideline and point out parallels to the associative case whenever possible. It turns out that this problem at hand results as a composite system. The Jordan algebraic approach has the particular advantage that, given the density matrix of the composite system, one can identify the state vectors corresponding to the subsystems. We accomplish this by defining a trace partially on the ambient space. Finally at the end, we are able to derive a Schrödinger-like time evolution equation that involves Jordan algebraic associators for this specific problem. This approach we think is something new with further possible applications.
2 Hilbert-Schmidt Operator Formulation
It turns out that given a pure state $\psi$ of a $C^{*}$-algebra $\mathcal{A}$, one can construct a Hilbert space $\mathscr{H}$, and write down a representation between $\mathcal{A}$ and the collection of bounded operators on $\overline{\mathcal{A}/\mathcal{I}_{\psi}}$, where $\mathcal{I}_{\psi}$ is defined as follows:
$\mathcal{I}_{\psi}=\{A\in\mathcal{A}:\psi(A^{*}A)=0\}.$
This construction lies at the heart of the celebrated the Gelfand-Naimark-Segal (GNS) theory. But this method is not suitable for the Jordan algebraic approach. Therefore, one demands another Hilbert space construction, which is based on traces of operators instead of their states.
From now on, let us assume that $\mathcal{A}$ is the algebra $\mathcal{B}(\mathscr{H})$ of all bounded operators on $\mathscr{H}$, which is well-known to be a von Neumann algebra. As a result of this, there exists a semi-finite, faithful, normal trace function on $\mathcal{A}$ [35]. We will then denote positive, trace class elements by $\mathcal{A}_{1}$, and those $A\in\mathcal{A}$ with $tr(A^{*}A)<\infty$ by $\mathcal{A}_{2}$. One can check that $\mathcal{A}_{2}$ is an ideal in $\mathcal{A}$. In particular, for all $A,B\in\mathcal{A}_{2}$, we have
$\langle A,B\rangle=trA^{*}B,\ \ \ \ \ \|A\|_{tr}=\sqrt{trA^{*}A}.$
Inspired by the GNS construction, one obtains the Hilbert space of Hilbert-Schmidt operators in the form
^$\hat{\mathscr{H}}=\overline{\mathcal{A}_{2}}^{tr}.$
Since $\mathcal{A}_{2}$ is an ideal, and $\mathcal{A}$ is associative, $\mathcal{A}$ can be represented over $\mathcal{A}_{2}$ by the multiplication operator which is defined by
$\pi(A)B=AB$
for all $A\in\mathcal{A}$, $B\in\mathcal{A}_{2}$. Thus, $\pi$ can be extended to a faithful representation of $\mathcal{A}$ on $\hat{\mathscr{H}}$ [36]. If we take a state $\psi$ represented by a density matrix $\rho\in\mathcal{A}_{1}$, then there is at least one $B\in\mathcal{A}_{2}$ such that $\rho=BB^{*}$ holds. At this point, it is worth mentioning that the expectation value of an observable $A\in\mathcal{A}$ is given by
$\psi(A)=tr\rho A=trBB^{*}A=trB^{*}AB=\langle B|A|B\rangle.$
Next let $U:\mathscr{H}\xrightarrow{}\mathscr{H}$ be a bounded, unitary operator. If we take any vector $B\in\mathcal{A}_{2}$, then $BU$ becomes a vector in $\hat{\mathscr{H}}$ describing the same vector as $B$ does. Therefore, the representation of states in $\hat{\mathscr{H}}$ is not unique, more precisely the states in $\hat{\mathscr{H}}$ are represented by right unitary orbits defined by
$BUU^{*}B^{*}=BB^{*}=\rho.$
On the other hand, every Hilbert-Schmidt operator $B\in\mathcal{A}_{2}$ with nonzero trace leads to a density matrix
$\rho_{B}:=\frac{BB^{*}}{tr(BB^{*})}.$
To emphasize the difference between $\mathscr{H}$ and $\hat{\mathscr{H}}$, note that now, every (pure and mixed) state is represented by at least one vector in $\hat{\mathscr{H}}$. It turns out that $\mathscr{H}$ and $\hat{\mathscr{H}}$ are connected in the following algebraic sense [36]
^=*¯$\hat{\mathscr{H}}=\overline{\mathscr{H}\otimes\mathscr{H}^{*}}^{tr}.$
We will now restrict this formalism to the physical system of quantum harmonic oscillator, which turns out to be very crucial in the formulation of the Landau problem. It is then necessary to focus on the application of Hilbert-Schmidt operators, which can be viewed as the collection of bounded operators acting on the configuration space
$\mathscr{H}:=span\left\{|n\rangle:\frac{1}{\sqrt{n!}}(a^{\dagger})^{n}|0\rangle\right\}_{n=0}^{\infty}.$
Furthermore, this space is isomorphic to the boson Fock space =$\mathcal{F}=\{|n\rangle\}_{n=0}^{\infty}$, where the annihilation and creation operators obey the Fock algebra $[a,a^{\dagger}]=1$. Observe that the physical states of the system represented on $\hat{\mathscr{H}}$, the set of Hilbert-Schmidt operators, in other words the set of all bounded operators on $\mathscr{H}=L^{2}(\mathbb{R})$ equipped with the scalar product
$\langle X|Y\rangle=tr(X^{*}Y)=\sum_{k}{\langle\phi_{k}|X^{*}Y\phi_{k}\rangle},$
where $\{\phi_{k}\}_{k=0}^{\infty}$ is an orthonormal basis of $\mathscr{H}$. Moreover, the basis vectors of $\hat{\mathscr{H}}$ are given by
$\phi_{nl}:=|\phi_{n}\rangle\langle\phi_{l}|,\ \ \ n,l=0,1,2,\dots\ .$
Here, we should note that $\mathscr{H}\otimes\mathscr{H}^{*}$ is a pre-Hilbert space containing all finite sums of the form
$\sum_{j,k=0}^{n}\lambda_{jk}|\phi_{j}\rangle\otimes|\phi_{k}\rangle^{*},$
where a basis of *¯$\overline{\mathscr{H}\otimes\mathscr{H}^{*}}^{tr}$ is
$\{|\phi_{j}\rangle\otimes|\phi_{k}\rangle^{*}\}_{j,k=0}^{\infty}.$
Setting $|\phi_{j}\rangle\otimes|\phi_{k}\rangle^{*}=|\phi_{j}\rangle\langle\phi_{k}|$, $\hat{\mathscr{H}}$ admits an orthonormal basis $\{\phi_{jk}\}_{j,k=0}^{\infty}$ such that
$\phi_{jk}:=|\phi_{j}\rangle\langle\phi_{k}|.$
Note that $\hat{\mathscr{H}}$ can also be described as the Hilbert space of square integrable functions on (,)($\infty$,$\infty$)( - ∞ , ∞ ) by
^=$\hat{\mathscr{H}}=\left\{\psi(x_{1},x_{2}):\psi(x_{1},x_{2})\in\mathcal{B}(\mathscr{H}),\ tr(\psi(x_{1},x_{2})^{\dagger}\psi(x_{1},x_{2}))<\infty\right\}.$ (1)
Recall that $\hat{\mathscr{H}}$ is also the set of bounded operators, described in (1), acting on the configuration space $\mathscr{H}$. Clearly, a general element in this space can be written as
$|\psi\rangle=\sum_{n,m=0}^{\infty}c_{n,m}|n,m\rangle,$
where $\{|n,m\rangle:=|n\rangle\langle m|\}_{n,m=0}^{\infty}$ is a basis of $\hat{\mathscr{H}}$ satisfying the inner product condition
$\langle n_{1},m_{1}|n_{2},m_{2}\rangle=tr[(|n_{1}\rangle\langle m_{1}|)^{\dagger}|n_{2}\rangle\langle m_{2}|]=\delta_{n_{1},n_{2}}\delta_{m_{1},m_{2}}.$
3 Jordan Algebraic Formulation
3.1 Hilbert Space Construction
In this section, we will present a Jordan algebraic GNS-type construction. But this time, our development will be based on the JB algebra of observables. The idea that a non-associative algebra of observables equipped with a tri-cyclic trace leading to a Hilbert space construction is very fruitful. For more details on this construction and its applications, we especially recommend the recent work [37]. It turns out that there is an analogue of von Neumann algebras in the context of JB algebras, which are called as JBW algebras [38]. One can work with normal, faithful, semi-finite traces on a JBW algebra $\mathcal{A}$ [39]. Trace on a Jordan algebra is defined as a weight and it satisfies the cyclicity condition
$tr(A\bullet(B\bullet C))=tr((A\bullet B)\bullet C).$
Also for positive elements $A,B\in\mathcal{A}$, and positive real numbers $\lambda$, we have
$tr(A+B)=trA+trB,$
$tr\lambda A=\lambda trA.$
Similarly, we will then denote positive, trace class elements by $\mathcal{A}_{1}$, and those $A\in\mathcal{A}$ with $tr(A^{2})<\infty$ by $\mathcal{A}_{2}$. One can check that $\mathcal{A}_{2}$ is a Jordan ideal in $\mathcal{A}$. Trace induces a bilinear, symmetric, real scalar product on $\mathcal{A}_{2}$, and in particular, we have
$\langle A|B\rangle=tr(A\bullet B),\ \ \ \ \ \|A\|_{tr}=\sqrt{trA^{2}}.$
Similar to the associative case, closure with respect to this norm yields a real Hilbert space
^$\hat{\mathscr{H}}_{J}=\overline{\mathcal{A}_{2}}^{tr}.$
Next we aim to write down a Jordan representation from $\mathcal{A}$ to this Hilbert space. Note that $\mathcal{A}_{2}$ is a Jordan module since for all $\omega\in\mathcal{A}_{2}$, and $A,B\in\mathcal{A}$, we have [38], [40]
$A\bullet\omega=\omega\bullet A,$
$A^{2}\bullet(A\bullet\omega)=A\bullet(A^{2}\bullet\omega),$
$2A\bullet(B\bullet(A\bullet\omega))+(B\bullet A^{2})\bullet\omega=2(A\bullet B)\bullet(A\bullet\omega)+A^{2}\bullet(B\bullet\omega).$
We may then define
$\pi_{J}(A):\mathcal{A}_{2}\xrightarrow{}\mathcal{A}_{2},\ \ \ \pi_{J}(A)B=A\bullet B$
for all $A\in\mathcal{A},B\in\mathcal{A}_{2}$. It is easy to check that $\pi_{J}$ satisfies the following properties
$\pi_{J}(A^{2})\bullet\pi_{J}(A)=\pi_{J}(A)\bullet\pi_{J}(A^{2}),$
$2\pi_{J}(A)\bullet\pi_{J}(B)\bullet\pi_{J}(A)+\pi_{J}(B\bullet A^{2})=2\pi_{J}(A\bullet B)\bullet\pi_{J}(A)+\pi_{J}(A^{2})\bullet\pi_{J}(B).$
Therefore, it becomes a Jordan representation. Moreover, due to the continuity of multiplication, $\pi_{J}$ can be extended to whole space ^$\hat{\mathscr{H}}_{J}$.
At this point, we should stress the relationship between ^$\hat{\mathscr{H}}_{J}$ and $\hat{\mathscr{H}}$. It turns out that ^$\hat{\mathscr{H}}_{J}$ is the real subspace of self-adjoint Hilbert-Schmidt operators embedded in $\hat{\mathscr{H}}$ [41]. Turning back to the problem of quantum harmonic oscillator, ^$\hat{\mathscr{H}}_{J}$ is a real span of following three family of operators
$|\phi_{j}\rangle\langle\phi_{j}|,$
$\frac{1}{2}(|\phi_{j}\rangle\langle\phi_{k}|+|\phi_{k}\rangle\langle\phi_{j}|),$ (2)
$\frac{i}{2}(|\phi_{j}\rangle\langle\phi_{k}|-|\phi_{k}\rangle\langle\phi_{j}|),$
where $j>k$, and $j,k$ ranges from 00 to $\infty$. Note that these elements form a basis for ^$\hat{\mathscr{H}}_{J}$, since they are self-adjoint. Here, $\{\phi_{k}\}_{k=0}^{\infty}$ is an orthonormal basis for $L^{2}(\mathbb{R})$ which can be written in terms of Hermite polynomials in the position space representation.
In a Jordan algebra, one can define the associator in the form
$[A,B,C]:=(A\bullet B)\bullet C-A\bullet(B\bullet C).$
Our aim is to find a connection between the commutator defined over the associative algebra $\mathcal{B}$ and the associator over $\mathcal{A}$, which is the self-adjoint part of $\mathcal{B}$.
Proposition 1.
Every self-adjoint element $H\in\mathcal{A}$ can be written as a finite sum of commutators in the form
$H=i\sum_{j=1}^{N}[{H_{L}}_{j},{H_{R}}_{j}],$ (3)
where ${H_{L}}_{j}$, and ${H_{R}}_{j}$ are self-adjoint operators in $\mathcal{A}$.
Proof.
Let $Aut^{0}(\mathcal{A})$ be the connected $1-$component of the Banach Lie group of all automorphisms of $\mathcal{A}$. Let $A$ be a self-adjoint element in $\mathcal{A}$. We define an operator $U$ as $U=\exp(iA)$, which is clearly a unitary operator. By [42], $Aut^{0}(\mathcal{A})$ is known to be algebraically generated by the symmetries in $\mathcal{A}$. Moreover by Theorem 1.6 of [42], each unitary operator can be written as a finite product of symmetries in $\mathcal{A}$. Combining these facts, we deduce that $U\in Aut^{0}(\mathcal{A})$. Armed with this, we can now apply Theorem 2.5 of [42] to obtain
$U=\exp\left(\sum_{j=1}^{n}[L_{a_{j}},L_{b_{j}}]\right),$
where $a_{j},b_{j}\in\mathcal{A}$ for all $j$, and $L_{a}b:=a\bullet b$ is the left multiplication operator. Consequently, we have
$\exp\left(iA-\sum_{j=1}^{n}[L_{a_{j}},L_{b_{j}}]\right)=1.$
Our next goal is to verify that
$\exp\left(t\left(iA-\sum_{j=1}^{n}[L_{a_{j}},L_{b_{j}}]\right)\right)$
is a one parameter family for $t\in\mathbb{R}$. First, this operator depends continuously on $t$. Second, when $t=0$, the operator reduces to the identity operator. Lastly, for any $t,s\in\mathbb{R}$, it follows from the properties of the exponential of an operator that
$\exp\left(\left(t+s\right)\left(iA-\sum_{j=1}^{n}[L_{a_{j}},L_{b_{j}}]\right)\right)=\exp\left(t\left(iA-\sum_{j=1}^{n}[L_{a_{j}},L_{b_{j}}]\right)+s\left(iA-\sum_{j=1}^{n}[L_{a_{j}},L_{b_{j}}]\right)\right)$
$=\exp\left(t\left(iA-\sum_{j=1}^{n}[L_{a_{j}},L_{b_{j}}]\right)\right)\exp\left(s\left(iA-\sum_{j=1}^{n}[L_{a_{j}},L_{b_{j}}]\right)\right).$
By the uniquness of generators of one-parameter families of operators (see Stone’s theorem [43]), we may now conclude that
$A=i\sum_{j=1}^{n}[L_{a_{j}},L_{b_{j}}].$
And (3) follows. ∎
3.2 Example: Non-commutative Landau Problem
[…]
$H_{\theta}=\frac{1}{2m}\left(p_{x}+\frac{eB}{2c}y\right)^{2}+\frac{1}{2m}\left(p_{y}-\frac{eB}{2c}x\right)^{2}+\frac{m\omega^{2}}{2}(x^{2}+y^{2}),$
where $\theta=-\frac{\hbar c}{eB}$ is the non-commutative parameter, $\omega_{L}=\frac{eB}{mc}$ is the Larmour frequency. Position and momentum operators satisfy the following nonzero commutation relations of the non-commutative Heisenberg algebra
$[x,y]=i\theta,\ \ \ [x,p_{x}]=[y,p_{y}]=i\hbar.$ (4)
Here, $x$ and $y$ denotes the center of the cyclotron motion, and they can be written in terms of commuting position coordinates $\tilde{x}$ and $\tilde{y}$ as follows
$x=\tilde{x}+\frac{c}{eB}\left(p_{y}-\frac{e}{c}A_{y}\right),\ \ y=\tilde{y}-\frac{c}{eB}\left(p_{x}-\frac{e}{c}A_{x}\right).$
Next, we define new operators in $\mathcal{B}$, corresponding to complex variables, related to chiral decomposition of non-commutative Landau problem. It is worth remarking that our approach here is reminiscent of the introduction of complex coordinates in the Landau problem pertaining to the Bloch waves quasi-periodicity on a torus (see [44]). Precisely, we have
$z=\frac{1}{\sqrt{2}}(x+iy),\ \ \ \overline{z}=\frac{1}{\sqrt{2}}(x-iy).$
Corresponding momentum operators are
$p_{z}=\frac{1}{\sqrt{2}}(p_{x}-ip_{y}),\ \ \ p_{\overline{z}}=\frac{1}{\sqrt{2}}(p_{x}+ip_{y}).$
These operators can be shown to satisfy the following commutation relations in $\mathcal{B}$
$[p_{z},z]=[p_{\overline{z}},\overline{z}]=-i\hbar,$
$[p_{z},\overline{z}]=[p_{\overline{z}},z]=[z,\overline{z}]=[p_{z},p_{\overline{z}}]=0.$
We set two pairs of creation and annihilation operators in $\mathcal{B}$
$A_{+}=\xi\frac{\overline{z}}{2}+\frac{i}{\xi\hbar}p_{z},\ \ \ A_{+}^{\dagger}=\xi\frac{z}{2}-\frac{i}{\xi\hbar}p_{{\overline{z}}},$
$A_{-}=\xi\frac{z}{2}+\frac{i}{\xi\hbar}p_{{\overline{z}}},\ \ \ A_{-}^{\dagger}=\xi\frac{\overline{z}}{2}-\frac{i}{\xi\hbar}p_{z},$
where
$\xi=\sqrt[4]{\frac{m^{2}\omega^{2}/\hbar^{2}+m^{2}{\omega_{L}}^{2}/4\hbar^{2}}{1-(m\omega_{L}\theta/2)+(m^{2}\omega^{2}\theta^{2}/16)+(m^{2}{\omega_{L}}^{2}\theta^{2}/64)}}.$
The nonzero commutation relations between these operators are
$[A_{+},A_{+}^{\dagger}]=1,\ [A_{-},A_{-}^{\dagger}]=1.$
Finally, defining $\Omega_{\pm}:=\Omega\pm\frac{\tilde{\omega}_{L}}{2}$, where
$\Omega=\sqrt{\omega^{2}+\frac{\omega_{L}^{2}}{4}-\frac{m\omega_{L}\omega^{2}\theta}{2}-\frac{m\omega_{L}^{3}\theta}{8}+\left(\omega^{2}+\frac{\omega_{L}^{2}}{4}\right)\left(\left(\frac{m\omega\theta}{4}\right)^{2}+\left(\frac{m\omega_{L}\theta}{8}\right)^{2}\right)},$
$\tilde{\omega}_{L}=\omega_{L}\left(1-\left(\frac{\omega_{L}}{4}+\frac{\omega^{2}}{\omega_{L}}\right)m\theta\right),$
one can write $H_{\theta}$ as follows [45], [33]
$H_{\theta}=\hbar\Omega_{+}\left(N_{+}+\frac{1}{2}\right)+\hbar\Omega_{-}\left(N_{-}+\frac{1}{2}\right),\ \ \ N_{\pm}=A_{\pm}^{\dagger}A_{\pm}.$ (5)
Let
$H_{\pm}=\hbar\Omega_{\pm}\left(N_{\pm}+\frac{1}{2}\right).$
We define the operator $H_{+}\otimes\mathbb{I}+\mathbb{I}\otimes H_{-}$ to be the infinitesimal generator of the strongly continuous one-parameter unitary group $e^{itH_{+}}\otimes e^{itH_{-}}$. Thus by Stone’s theorem, $H_{+}\otimes\mathbb{I}+\mathbb{I}\otimes H_{-}$ is self-adjoint (for details of the above argument see [33] and page 432, Definition 19.15 of [46]). Using that, (5) can be written as
$H_{\theta}=H_{+}\otimes\mathbb{I}_{\hat{\mathscr{H}}_{J,-}}+\mathbb{I}_{\hat{\mathscr{H}}_{J,+}}\otimes H_{-}.$
Here ^$\hat{\mathscr{H}}_{J,+}\otimes\hat{\mathscr{H}}_{J,-}$ is the real Jordan-Hilbert space for this problem described in (2), which is the domain of $H_{\theta}$, and $\mathbb{I}_{\hat{\mathscr{H}}_{J,\pm}}$ are the corresponding identity operators.
Theorem 1.
Nonzero commutation relations in $\mathcal{B}$, given in $(4)$, can be written as associator relations in $\mathcal{A}$ as follows
$[y_{L},x,y_{R}]=\theta,\ \ \ [{p_{x}}_{L},x,{p_{x}}_{R}]=\hbar,\ \ \ [{p_{y}}_{L},y,{p_{y}}_{R}]=\hbar,$ (6)
where
$y_{L}={p_{x}}_{L}={p_{y}}_{L}=4H_{\theta},$
$y_{R}=\left(\frac{1}{2\hbar\Omega_{-}}-\frac{1}{2\hbar\Omega_{+}}\right)x-\left(\frac{1}{\xi^{2}\hbar^{2}\Omega_{+}}+\frac{1}{\xi^{2}\hbar^{2}\Omega_{-}}\right)p_{y},$
${p_{x}}_{R}=\left(\frac{\xi^{2}}{4\Omega_{+}}+\frac{\xi^{2}}{4\Omega_{-}}\right)x+\left(\frac{1}{2\hbar\Omega_{+}}-\frac{1}{2\hbar^{2}\Omega_{-}}\right)p_{y},$
${p_{y}}_{R}=\left(\frac{\xi^{2}\hbar}{4\Omega_{+}}+\frac{\xi^{2}\hbar}{4\Omega_{-}}\right)y+\left(\frac{1}{2\Omega_{-}}-\frac{1}{2\Omega_{+}}\right)p_{x}.$
Proof.
In the case of a special Jordan algebra, an associator can be expressed as a double commutator
$[A,B,C]=\frac{1}{4}[B,[A,C]].$ (7)
In order to find the operators explicitly, one needs to write $x,y,p_{x}$, and $p_{y}$ in terms of newly defined creation and annihilation operators. It turns out that
$x=\frac{1}{\sqrt{2}\xi}(A_{-}+A_{+}+A_{+}^{\dagger}+A_{-}^{\dagger}),$
$y=\frac{i}{\sqrt{2}\xi}(-A_{-}+A_{+}-A_{+}^{\dagger}+A_{-}^{\dagger}),$
$p_{x}=\frac{i\xi\hbar}{2\sqrt{2}}(-A_{-}-A_{+}+A_{+}^{\dagger}+A_{-}^{\dagger}),$
$p_{y}=\frac{\xi\hbar}{2\sqrt{2}}(-A_{-}+A_{+}+A_{+}^{\dagger}-A_{-}^{\dagger}).$
The following commutation relation can be written for $y$
$\left[\frac{H_{\theta}}{\Omega_{+}},A_{+}+A_{+}^{\dagger}\right]+\left[\frac{H_{\theta}}{\Omega_{-}},-A_{-}-A_{-}^{\dagger}\right]$
$=\hbar([A_{+}^{\dagger}A_{+},A_{+}]+[A_{+}^{\dagger}A_{+},A_{+}^{\dagger}]+[A_{-}^{\dagger}A_{-},-A_{-}]+[A_{-}^{\dagger}A_{-},-A_{-}^{\dagger}])$
$=\hbar(-A_{+}+A_{+}^{\dagger}+A_{-}-A_{-}^{\dagger})=i\sqrt{2}\xi\hbar y.$
So we have,
$i\left[H_{\theta},\left(\frac{1}{2\hbar\Omega_{-}}-\frac{1}{2\hbar\Omega_{+}}\right)x-\left(\frac{1}{\xi^{2}\hbar^{2}\Omega_{+}}+\frac{1}{\xi^{2}\hbar^{2}\Omega_{-}}\right)p_{y}\right]=y.$
Similarly, one can check that
$i\left[H_{\theta},\left(\frac{\xi^{2}}{4\Omega_{+}}+\frac{\xi^{2}}{4\Omega_{-}}\right)x+\left(\frac{1}{2\hbar\Omega_{+}}-\frac{1}{2\hbar^{2}\Omega_{-}}\right)p_{y}\right]=p_{x},$
$i\left[H_{\theta},\left(\frac{\xi^{2}\hbar}{4\Omega_{+}}+\frac{\xi^{2}\hbar}{4\Omega_{-}}\right)y+\left(\frac{1}{2\Omega_{-}}-\frac{1}{2\Omega_{+}}\right)p_{x}\right]=p_{y}.$
Using (7) and the commutation relation provided for $y$, $p_{x}$ and $p_{y}$, associator relations (6) follow. ∎
3.3 Density Matrices in the JBW-algebra Setting
It turns out that one needs a notion of density matrix in the JBW-algebra $\mathcal{A}$ in order to describe the state of a quantum system. For the purpose of applications to JBW-algebras, let us recall the following result from [47], [48].
Theorem.
Let $\mathcal{A}$ be a commutative, order unit algebra that is the dual of a base norm space $V$ such that the multiplication in $\mathcal{A}$ is separately $w^{*}$- continuous. Then for each $a\in\mathcal{A}$ and each $\epsilon>0$ there are orthogonal projections $p_{1},p_{2},\dots,p_{n}$ in the $w^{*}$-closed subalgebra generated by $a$ and $1$, and scalars $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$ such that
$\|a-\sum_{i=1}^{n}\lambda_{i}p_{i}\|<\epsilon.$
The set of finite linear combinations of orthogonal projections is norm dense in $\mathcal{A}^{**}$, here $\mathcal{A}^{**}$ denotes the bidual of $\mathcal{A}$, which is a JBW-algebra.
We will assume in Theorem 3 that any element can be written as a finite linear combination of orthonormal projection operators. In particular, for any $B\in\hat{\mathscr{H}}_{J}$ we have
$B=\sum_{i=1}^{n}\lambda_{i}P_{i},$ (8)
where $\lambda_{i}$’s are real scalars, and $P_{i}$’s denoting the orthonormal operators. And for a general element, we use the fact that it can be approximated in norm by linear combinations of orthonormal projections. So, the general case follows from the norm continuity of multiplication.
Definition 1.
A self-adjoint operator $\rho\in\mathcal{A}$ is a density matrix if $\rho$ is non-negative and $tr(\rho)=1$.
Theorem 2.
Assume that $\rho$ is a density matrix on ^$\hat{\mathscr{H}}_{J}$, then the map $\Phi_{\rho}:\mathcal{A}\xrightarrow{}\mathbb{R}$ given by
$\Phi_{\rho}(A)=tr(\rho\bullet A)$
is a family of expectation values.
Proof.
In order to show this, we need to verify the following three properties of $\Phi$.
(i) $\Phi_{\rho}(\mathbb{I})=1$,
(ii) $\Phi_{\rho}(A)$ is real for all $A\in\mathcal{A}$, and $\Phi_{\rho}(A)$ is non-negative whenever $A$ is non-negative.
(iii) For any sequence $A_{n}\in\mathcal{A}$, if
$\lim_{n\rightarrow\infty}\|A_{n}\bullet\psi-A\bullet\psi\|=0$
holds for all $\psi\in\hat{\mathscr{H}}_{J}$, then $\Phi_{\rho}(A_{n})$ converges to $\Phi_{\rho}(A)$.
Note that (i) is immediate from the definition of a density matrix. Since $A\in\mathcal{A}$ is self-adjoint, $tr(\rho\bullet A)$ is real by definition. Let $B\in\hat{\mathscr{H}}_{J}$ be the non-negative, self-adjoint square root of $\rho$. It follows that $tr(A\bullet(B\bullet B))=tr(B\bullet(A\bullet B))\geq 0$, since $B\bullet(A\bullet B)\in\hat{\mathscr{H}}_{J}$ is self-adjoint and non-negative. This verifies condition (ii). Finally, assume $A_{n}\bullet\psi$ converges in norm to $A\bullet\psi$ for all $\psi\in\hat{\mathscr{H}}_{J}$. Then, $\|A_{n}\bullet\psi\|$ is bounded for each fixed $\psi$. By the principle of uniform boundedness, there is a constant $C$ such that $\|A_{n}\|\leq C$. Let $P_{i}$’s denote the orthonormal projection operators introduced in (8), we have
$|\langle P_{i},(B\bullet(A_{n}\bullet B))\bullet P_{i}\rangle|=|\langle P_{i},(A_{n}\bullet(B\bullet P_{i}))\bullet B\rangle|$
$=|\langle(B\bullet P_{i}),A_{n}\bullet(B\bullet P_{i}))\rangle|\leq C\|B\bullet P_{i}\|^{2},$
$\lim_{m\rightarrow\infty}\sum_{i=1}^{m}\|B\bullet P_{i}\|^{2}=\lim_{m\rightarrow\infty}\sum_{i=1}^{m}\langle B\bullet P_{i},B\bullet P_{i}\rangle=\lim_{m\rightarrow\infty}\sum_{i=1}^{m}\langle P_{i},\rho\bullet P_{i}\rangle=tr(\rho)<\infty.$
Since $A_{n}\bullet(B\bullet P_{i})$ converges to $A\bullet(B\bullet P_{i})$ for each $i$, by dominated convergence
$tr(B^{2}\bullet A)=tr(B\bullet(A\bullet B))=\lim_{m\rightarrow\infty}\sum_{i=1}^{m}\langle P_{i},(B\bullet(A\bullet B))\bullet P_{i}\rangle$
$=\lim_{m\rightarrow\infty}\lim_{n\rightarrow\infty}\sum_{i=1}^{m}\langle P_{i},(B\bullet(A_{n}\bullet B))\bullet P_{i}\rangle$
$=\lim_{m\rightarrow\infty}\sum_{i=1}^{m}\langle P_{i},(B\bullet(A\bullet B))\bullet P_{i}\rangle=\lim_{n\rightarrow\infty}tr(B\bullet(A_{n}\bullet B))=\lim_{n\rightarrow\infty}tr(B^{2}\bullet A_{n}).$
For a normal state $B\in\hat{\mathscr{H}}_{J}$, the expectation value of an observable $A\in\mathcal{A}$ is given by
$\Phi_{B}(A)=\langle B,A\bullet B\rangle=tr(B\bullet(A\bullet B))=tr(B^{2}\bullet A)=tr(\rho\bullet A),$
where $\rho=B^{2}$ is the corresponding density matrix.
Definition 2.
A density matrix is called a pure state if it can be written as a projection operator in $\mathcal{A}$, otherwise it is called a mixed state.
Now, we can revisit the non-commutative Landau problem, where the Hilbert space for the composite system is the Hilbert tensor product ^$\hat{\mathscr{H}}_{J,+}\otimes\hat{\mathscr{H}}_{J,-}$ of the Hilbert spaces ^$\hat{\mathscr{H}}_{J,+}$ and ^$\hat{\mathscr{H}}_{J,-}$ describing the subsystems. Assuming $\rho$ is a density matrix on ^$\hat{\mathscr{H}}_{J,+}\otimes\hat{\mathscr{H}}_{J,-}$, there exists a unique density matrix $\rho_{+}$ on ^$\hat{\mathscr{H}}_{J,+}$ with the property that $tr(\rho_{+}\bullet A)=tr(\rho\bullet(A\otimes\mathbb{I}_{\hat{\mathscr{H}}_{J,-}}))$ for all $A\in\mathcal{B}(\hat{\mathscr{H}}_{J,+})$. Similarly, there exists a unique density matrix $\rho_{-}$ on ^$\hat{\mathscr{H}}_{J,-}$ with the property that $tr(\rho_{-}\bullet B)=tr(\rho\bullet(\mathbb{I}_{\hat{\mathscr{H}}_{J,+}}\otimes B))$ for all $B\in\mathcal{B}(\hat{\mathscr{H}}_{J,-})$. Since, the state of the first system is independent of the state of the second system, the density matrix $\rho$ is of the form $\rho=\rho_{+}\otimes\rho_{-}$ for this problem.
Theorem 3.
A pure state is being represented in ^$\hat{\mathscr{H}}_{J,+}\otimes\hat{\mathscr{H}}_{J,-}$ by ±$\pm P_{0,+}\otimes\pm P_{0,-}$, where $P_{0,+}$ and $P_{0,-}$ are idempotent elements on ^$\hat{\mathscr{H}}_{J,+}$ and ^$\hat{\mathscr{H}}_{J,-}$, respectively.
Proof.
To prove our claim, we consider a pure state $\rho=\rho_{+}\otimes\rho_{-}$, given by an idempotent $P_{0,+}\otimes P_{0,-}$. Without loss of generality, we can take $P_{0,+}$ as the first basis vector of ^$\hat{\mathscr{H}}_{J,+}$, and $P_{0,-}$ as the first basis vector of ^$\hat{\mathscr{H}}_{J,-}$. We denote the other basis vectors as $P_{i,\pm}$. Let $B$ be the square root of $\rho$. Then we have
$B=\left(\sum_{i=0}^{n}\lambda_{i}P_{i,+}\right)\otimes\left(\sum_{i=0}^{n}\mu_{i}P_{i,-}\right).$
Using this, $\rho$ can be written as
$\rho=B^{2}=\left(\sum_{i=0}^{n}\lambda_{i}^{2}P_{i,+}\right)\otimes\left(\sum_{i=0}^{n}\mu_{i}^{2}P_{i,-}\right),$
Taking traces of both sides, we obtain
$\left(\sum\lambda_{i}^{2}\right)\left(\sum\mu_{i}^{2}\right)=1.$
Finally, we have
$P_{0,+}\otimes P_{0,-}=\left(\sum_{i=0}^{n}\lambda_{i}^{2}P_{i,+}\right)\otimes\left(\sum_{i=0}^{n}\mu_{i}^{2}P_{i,-}\right).$
The above formula clearly implies that $\lambda_{0}=\pm 1$, $\mu_{0}=\pm 1$, and $P_{i,\pm}=0$ for $i>0$. ∎
Now, we will define new observables before stating our next theorem giving the time evolution equation for the state vectors of non-commutative Landau problem. Let
$X_{+}=C_{1_{+}}(A_{+}^{\dagger}+A_{+}),\ \ \ \ \ X_{-}=C_{1_{-}}(A_{-}^{\dagger}+A_{-}),$
be the position operators corresponding to subsystems. And
$P_{+}=iC_{2_{+}}(A_{+}^{\dagger}-A_{+}),\ \ \ \ \ P_{-}=iC_{2_{-}}(A_{-}^{\dagger}-A_{-}),$
be the corresponding momentum operators, where
$C_{1_{+}}=\sqrt{\frac{\hbar}{2m\Omega_{+}}},\ \ \ C_{2_{+}}=\sqrt{\frac{\hbar m\Omega_{+}}{2}},\ \ \ C_{1_{-}}=\sqrt{\frac{\hbar}{2m\Omega_{-}}},\ \ \ C_{2_{-}}=\sqrt{\frac{\hbar m\Omega_{-}}{2}}.$
Theorem 4.
Given a state vector $v_{+}\otimes v_{-}\in\hat{\mathscr{H}}_{J,+}\otimes\hat{\mathscr{H}}_{J,-}$, the time evolution equation can be written as follows
$\frac{d}{dt}(v_{+}\otimes v_{-})=\frac{-4}{\hbar}((([L_{S_{1_{+}}},L_{R_{1_{+}}}]+[L_{S_{2_{+}}},L_{R_{2_{+}}}])v_{+}\otimes v_{-})+(v_{+}\otimes([L_{S_{1_{-}}},L_{R_{1_{-}}}]+[L_{S_{2_{-}}},L_{R_{2_{-}}}])v_{-})),$
where $H_{+}=i[R_{1_{+}},S_{1_{+}}]+i[R_{2_{+}},S_{2_{+}}]$, $H_{-}=i[R_{1_{-}},S_{2_{-}}]+i[R_{2_{-}},S_{2_{-}}]$,
$R_{1_{+}}=-\frac{\Omega_{+}}{4}(X_{+}A_{+}+A_{+}^{\dagger}X_{+}),$
$R_{2_{+}}=-\frac{\Omega_{+}}{4}(2C_{1_{+}}P_{+}A_{+}+2C_{1_{+}}A_{+}^{\dagger}P_{+}),$
$S_{1_{+}}=P_{+}A_{+}+A_{+}^{\dagger}P_{+},$
$S_{2_{+}}=2C_{1_{+}}A_{+}^{\dagger}A_{+},$
$R_{1_{-}}=-\frac{\Omega_{-}}{4}(X_{-}A_{-}+A_{-}^{\dagger}X_{-}),$
$R_{2_{-}}=-\frac{\Omega_{-}}{4}(2C_{1_{-}}P_{-}A_{-}+2C_{1_{-}}A_{-}^{\dagger}P_{-}),$
$S_{1_{-}}=P_{-}A_{-}+A_{-}^{\dagger}P_{-},$
$S_{2_{-}}=2C_{1_{-}}A_{-}^{\dagger}A_{-},$
Proof.
Let us first see that $H_{+}=i[R_{1_{+}},S_{1_{+}}]+i[R_{2_{+}},S_{2_{+}}]$, $H_{-}=i[R_{1_{-}},S_{2_{-}}]+i[R_{2_{-}},S_{2_{-}}]$. We will only show the first equality, since two subsystems are independent identical copies. Note that
$\frac{-i}{C_{1_{+}}C_{2_{+}}}[X_{+}A_{+}+A_{+}^{\dagger}X_{+},P_{+}A_{+}+A_{+}^{\dagger}P_{+}]$ $=[2A_{+}^{\dagger}A_{+}+A_{+}^{2}+{A_{+}^{\dagger}}^{2},{A_{+}^{\dagger}}^{2}-A_{+}^{2}]$
$=[2A_{+}^{\dagger}A_{+},{A_{+}^{\dagger}}^{2}]-[2A_{+}^{\dagger}A_{+},A_{+}^{2}]+2[A_{+}^{2},{A_{+}^{\dagger}}^{2}]$
$=2(A_{+}^{\dagger}[A_{+},{A_{+}^{\dagger}}^{2}]-[A_{+}^{\dagger},A_{+}^{2}]A_{+}+[A_{+},{A_{+}^{\dagger}}^{2}]A_{+}+A_{+}[A_{+},{A_{+}^{\dagger}}^{2}])$
$=4{A_{+}^{\dagger}}^{2}+4A_{+}^{2}+8A_{+}^{\dagger}A+4\mathbb{I}.$
Similarly,
$\frac{-i}{C_{2_{+}}}[P_{+}A_{+}+A_{+}^{\dagger}P_{+},A_{+}^{\dagger}A]$ $=[{A_{+}^{\dagger}}^{2},A_{+}^{\dagger}A_{+}]-[A_{+}^{2},A_{+}^{\dagger}A_{+}]$
$=-[A_{+}^{\dagger}A_{+},A_{+}^{\dagger}]A_{+}^{\dagger}-A_{+}^{\dagger}[A_{+}^{\dagger}A_{+},A_{+}^{\dagger}]+[A_{+}^{\dagger}A_{+},A_{+}]A_{+}+A_{+}[A_{+}^{\dagger}A_{+},A_{+}]$
$=-2{A_{+}^{\dagger}}^{2}-2A_{+}^{2}.$
Combining these, we get the required result. The time evolution equation for density matrices in a $C^{*}$-algebra is given by
$\dot{\rho}=-\frac{i}{\hbar}[H,\rho],$ (9)
where $H$ denotes the Hamiltonian. In the case of Jordan algebras, inner derivations are given in terms of the associator, which acts as a derivation on the second argument [38]. Assuming $H=i[R,S]$, where $R$ and $S$ are self-adjoint, (9) transforms into the following equation
$\dot{\rho}=\frac{-4}{\hbar}[R,\rho,S]=\frac{-4}{\hbar}[L_{S},L_{R}]\rho,$
where $L_{S}$ and $L_{R}$ are left multiplication operators with $S$ and $R$, respectively. This amounts to the Jordan-algebraic evolution equation for the density matrices. All density matrices can be written in the form $\rho=B^{2}$, for some $B\in\mathcal{A}_{2}$. Taking the derivative of both sides
$\dot{\rho}=2B\bullet\dot{B}.$
On the other hand, we also have
$\dot{\rho}=\frac{-4}{\hbar}[R,\rho,S]=\frac{-4}{\hbar}[R,B^{2},S]=\frac{-4}{\hbar}2B\bullet[R,B,S].$
Comparing them, one obtains
$\dot{B}=\frac{-4}{\hbar}[R,B,S]=\frac{-4}{\hbar}[L_{S},L_{R}]B.$
Let us remark that this equation can be extended to whole space ^$\hat{\mathscr{H}}_{J}$ by
$\dot{v}=\frac{-4}{\hbar}[R,v,S]=\frac{-4}{\hbar}[L_{S},L_{R}]v.$
Next, let $v_{+}\otimes v_{-}$ be an element of ^$\hat{\mathscr{H}}_{J,+}\otimes\hat{\mathscr{H}}_{J,-}$ corresponding to the density matrix $\rho_{+}\otimes\rho_{-}$, and let $H_{\theta}=H_{+}\otimes\mathbb{I}_{\hat{\mathscr{H}}_{J,-}}+\mathbb{I}_{\hat{\mathscr{H}}_{J,+}}\otimes H_{-}$ be the Hamiltonian, where $H_{+}=i[R_{1_{+}},S_{1_{+}}]+i[R_{2_{+}},S_{2_{+}}]$, $H_{-}=i[R_{1_{-}},S_{2_{-}}]+i[R_{2_{-}},S_{2_{-}}]$. Based on these, we may deduce that
$\frac{d}{dt}(\rho_{+}\otimes\rho_{-})=-\frac{i}{\hbar}[H_{+}\otimes\mathbb{I}_{\hat{\mathscr{H}}_{J,-}}+\mathbb{I}_{\hat{\mathscr{H}}_{J,+}}\otimes H_{-},\rho_{+}\otimes\rho_{-}]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =-\frac{i}{\hbar}([H_{+}\otimes\mathbb{I}_{\hat{\mathscr{H}}_{J,-}},\rho_{+}\otimes\rho_{-}]+[\mathbb{I}_{\hat{\mathscr{H}}_{J,+}}\otimes H_{-},\rho_{+}\otimes\rho_{-}])$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =-\frac{i}{\hbar}(([H_{+},\rho_{+}]\otimes\rho_{-})+(\rho_{+}\otimes[H_{-},\rho_{-}])).$
Using the discussion above, we may replace $-i[H,\rho]$ with $-4[L_{S_{1}},L_{R_{1}}]v-4[L_{S_{2}},L_{R_{2}}]v$ to arrive at the formula
$\frac{d}{dt}(v_{+}\otimes v_{-})=\frac{-4}{\hbar}((([L_{S_{1_{+}}},L_{R_{1_{+}}}]+[L_{S_{2_{+}}},L_{R_{2_{+}}}])v_{+}\otimes v_{-})+(v_{+}\otimes([L_{S_{1_{-}}},L_{R_{1_{-}}}]+[L_{S_{2_{-}}},L_{R_{2_{-}}}])v_{-})).$
Finally, the solution to this Jordan-Schrödinger equation can be written as
$v_{+}(t)\otimes v_{-}(t)=e^{\frac{-4t}{\hbar}([L_{S_{1_{+}}},L_{R_{1_{+}}}]+[L_{S_{2_{+}}},L_{R_{2_{+}}}])}v_{+}(0)\otimes e^{\frac{-4t}{\hbar}([L_{S_{1_{-}}},L_{R_{1_{-}}}]+[L_{S_{2_{-}}},L_{R_{2_{-}}}])}v_{-}(0).$
4 Concluding Remarks
In this work, for a two dimensional non-commutative space, a Jordan algebraic formulation of the Hamiltonian associated to the motion of an electron under a uniform magnetic field with a confining harmonic potential has been provided. The main result is given in Theorem 1, and the non-commutative parameter $\theta$ is written in terms of an associator in the Jordan algebra of observables. Secondly, a general understanding for the expectation value of an observable is provided in terms of the Jordan algebraic setting. In addition, the pure states are characterized for this specific problem. Lastly, a Jordan algebraic time evolution equation for the state vectors in the constructed Hilbert space and its solution are presented.
5 Acknowledgement
T.D. thanks The Turkish Academy of Sciences (TÜBA) for partial support.
6 Data Availability
Data sharing is not applicable to this article.